Assume that​ women's heights are normally distributed with a mean given by μ=64.9 in​, and a standard deviation given by σ=2.3 in. Complete parts a and b. a. If 1 woman is randomly​ selected, find the probability that her height is between 64.6 in and 65.6 in.

The probability is approximately _____. ​(Round to four decimal places as​ needed.)
Answer #1

Solution:

Given: The women's heights are normally distributed with a mean μ = 64.9 in​, and a standard deviation σ = 2.3 in.

Part a)   If 1 woman is randomly​selected, find the probability that her height is between 64.6 in and 65.6 in.

P( 64.6 < X < 65.6 ) =..............?

Find z score for x = 64.6 and for x = 65.6

1 -

-0.3 64.6-64.9 23 -=-0.13 23

65.6-64.9 23 = 0.7 23 。 = 0.30

Thus we get:

P( 64.6 < X < 65.6 ) = P( -0.13 < Z < 0.30 )

P( 64.6 < X < 65.6 ) = P( Z < 0.30 ) - P( Z < -0.13)

Look in z table for z = 0.3 and 0.00 as well as for  z = -0.1 and 0.03  and find corresponding area.
| 0.0 400 10.1 5998 | 0.2 $3 0.3-6179 .6554 10.5 6915 .01 .02 .03 15040 5080 5120 ,543854785517 ,5832 .5871 .5910 .6217 .6255

P( Z < 0.30) = 0.6179

-27 3 .00 .01 .02. 1.03 .04 .05 .06 .07 .08 .09 ] 1-3.4 .0003 .0003 .0003 10: 100 .000 .0003 .0003 .0003 .0002 | |-3.3 ,0005

P( Z < -0.13) = 0.4483

Thus

P( 64.6 < X < 65.6 ) = P( Z < 0.30 ) - P( Z < -0.13)

P( 64.6 < X < 65.6 ) = 0.6179 - 0.4483

P( 64.6 < X < 65.6 ) = 0.1696